OpenStudy (anonymous):

Given the system of equations presented here: 2x + 4y = 14 4x + y = 20 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated? A. Multiply the second equation by −4 to get −16x − 4y = −80 B.Multiply the second equation by −1 to get −4x − y = −20 C.Multiply the first equation by 2 to get 4x + 8y = 28 D.Multiply the first equation by −1 to get −2x − 4y = −14

Mehek (mehek14):

Multiply the first equation by 2 so when you subtract 4x - 4x x is eliminated

OpenStudy (anonymous):

ok Can you show me step by step i am really confused.

Mehek (mehek14):

when you multiply 2x + 4y = 14 by 2 you get 4x + 8y = 28

Mehek (mehek14):

now you have 4x + 8y = 28 4x + y = 20

OpenStudy (anonymous):